# QIANets: Quantum-Integrated Adaptive Networks for Reduced Latency and Improved Inference Times in CNN Models

### v1

@@ 4.3.3 ResNet-18
Lastly, ResNet-18 is a CNN characterized by its unique residual learning framework and shortcut connections that facilitate training in deep networks. Within our experiments, the QIANets framework targets the residual blocks in the model, reducing less significant weights and channels, detected by ECA’s straightforward 1D convolution. Feature maps are then combined by Multi-Scale Fusion and refined highlighting essential features across different scales.
@@
Loss Reduction: The loss reduction across trials also followed a clear downward trend. In Trial 3, with a layer sparsity of 0.4805 and a rank of 10, the validation loss dropped from 1.9847 in the first epoch to 0.6321 by the tenth epoch, indicating better model convergence. One of the standout results of this experiment was the final reduction in inference time by 36.4%, dropping to 0.00007 seconds/image, marking a improvement compared to the baseline ResNet’s 0.00011 seconds/image.
5 Conclusion
In this paper, we introduced the QIANets framework and applied it to three prominent convolutional neural networks—DenseNet, GoogLeNet, and ResNet—with the objective of reducing latency and improving inference times while maintaining minimal accuracy loss. Our experimental results showed varying degrees of success when applying quantum-inspired techniques, revealing important insights into the trade-offs between latency and accuracy across different architectures.
6 Limitations
While our results demonstrate the potential of QIANets and quantum-inspired principles in model compression, they also reflect on several factors that influence the performance of our approach:
  1. 1.
    Data Constraints: The evaluation was restricted to the relatively simple CIFAR-10 dataset, which may represent the full diversity, complexity, or scalability challenges encountered in real-world scenarios that larger datasets may exhibit. Additionally, due to the computational expense of each run, the approach was only tested on a limited number of trials.
  2. 2.
    Model Adaptation: The lack of adaptation across different model architectures may hinder the QIANets framework’s ability to optimize the balance between latency and accuracy. Performance in certain cases does not guarantee similar results across architectures without mandatory model-specific adjustments, which can complicate future adaptations.
  3. 3.
    Hardware Limitations: This study does not consider hardware-specific limitations. Our techniques have not yet been optimized for specialized hardware, such as custom FPGAs or GPUs, which could potentially reduce latency and improve data throughput.
  4. 4.
    Scope of Focus: The narrow focus on CNNs excludes newer, more advanced architectures, restricting the scope of our framework’s potential. Integrating our concept with transformers and other attention-based models could significantly expand QIANets’ applicability.
Future works should address these limitations by conducting more in-depth experiments that assess both the scalability and practical relevance of the quantum-inspired techniques.
Acknowledgments and Disclosure of Funding
Use unnumbered first level headings for the acknowledgments. All acknowledgments go at the end of the paper before the list of references. Moreover, you are required to declare funding (financial activities supporting the submitted work) and competing interests (related financial activities outside the submitted work). More information about this disclosure can be found at: https://neurips.cc/Conferences/2024/PaperInformation/FundingDisclosure.
Do not include this section in the anonymized submission, only in the final paper. You can use the ack environment provided in the style file to automatically hide this section in the anonymized submission.
@@ 18. Answer: [Yes]
  19. Justification: The code and instructions on how to reproduce the experiments are included in the supplementary material. Within the paper, the model is described in detail with code snippets and is reproducible using the information provided.
@@ 23. Answer: [Yes]
  24. Justification: Code is provided with instructions as supplemental material.

### v2

@@ 4.3.3 ResNet-18
Lastly, we tested on ResNet-18, a CNN characterized by its unique residual learning framework and shortcut connections. The QIANets framework targets the residual blocks in the model, reducing less significant weights and channels, detected by ECA’s straightforward 1D convolution. See Table 4.
Table 4: Training and validation metrics for ResNet-18
Metric Quantum-Inspired ResNet-18 Baseline ResNet-18
Training Loss 0.4078 0.6501
Validation Accuracy 90.25% 91.30%
Test Loss 0.6447 0.3195
Test Accuracy 87.11% 93.11%
Average Inference Time/Image 0.00007 seconds (36.4% faster) 0.00011 seconds
4.4 Analysis of the QIANets Framework
The QIANets framework achieves compression ratios of 1.9× for GoogleNet, 1.8× for DenseNet, and 1.6× for ResNet-18, demonstrating effective latency reductions. Each model showed consistent loss reduction, approached baseline accuracy post-fine-tuning, and achieved faster inference. Although results are slightly below some CNN compression methods [23], QIANets shows promise for quantum-inspired compression.
5 Conclusion
In this paper, we introduced the QIANets framework, applying it to DenseNet, GoogLeNet, and ResNet-18 to reduce latency and improve inference time while preserving accuracy. Our results highlight the potential of quantum-inspired techniques for CNN compression, yielding valuable insights into the trade-offs between latency and accuracy across various architectures.
6 Limitations
While our results demonstrate the potential of QIANets and quantum-inspired principles in model compression, they also highlight several factors influencing performance. Future work should address these limitations through more in-depth experiments assessing the scalability and practical relevance of quantum-inspired techniques.
  1. 1.
    Data Constraints: The evaluation was limited to the relatively simple CIFAR-10 dataset, which may not fully capture the diversity, complexity, or scalability challenges present in larger real-world datasets. Additionally, due to the computational expense, the approach was tested on a restricted number of trials.
  2. 2.
    Model Adaptation: The lack of adaptation across different architectures may hinder the QIANets framework’s ability to balance latency and accuracy. Performance in certain scenarios does not guarantee similar results across architectures without model-specific adjustments, complicating future adaptations.
  3. 3.
    Hardware Limitations: This study does not address hardware-specific limitations. Our techniques have yet to be optimized for specialized hardware, such as custom FPGAs or GPUs, which could further reduce latency and improve data throughput.
Acknowledgments and Disclosure of Funding
This work was completed through the Algoverse program, and we acknowledge the team for their support. We also thank Sean O’Brien and the anonymous reviewers for their valuable feedback.
Appendix A Extended Results and Model Analysis
A.1 GoogLeNet:
Performance Progression Across Epochs: During Trial 0 of optimization, the model began with a validation accuracy of 21.08% at Epoch 1 and quickly progressed to 70.18% by Epoch 10, indicating rapid learning during the first stages of training. The highest validation accuracy was 80.19% throughout Trial 5. The final quantum-inspired GoogLeNet’s test accuracy was 86.65%, a notable improvement from the earlier 80.19%, approaching the baseline accuracy of 94.29% after fine-tuning.
Loss Reduction: Over the 50 epochs, the model’s loss steadily decreased, showing consistent improvement. It began at 2.5205 in Epoch 1 trial 0 and reduced to 0.8256 by Epoch 50, effectively minimizing error throughout training. A key outcome of this experiment was the final 13.65% reduction in inference time, minimizing to 0.000835 seconds per image, which underscores the efficiency of our approach compared to the baseline GoogLeNet’s 0.000967 seconds. See Table 2.
A.2 DenseNet:
Performance Progression Across Epochs: The model began with a validation accuracy of 9.66% at Epoch 1 (Trial 0) but exhibited minimal improvement by Epoch 10, reaching 10.34%. However, Trial 1 demonstrated significant progress, starting at 27.53% and achieving a remarkable 81.33% by Epoch 10. The model achieved its highest validation accuracy of 86.65% in Trial 1, with a layer sparsity of 0.3779 (62% of weights pruned). After fine-tuning, the quantum-inspired DenseNet achieved a test accuracy of 88.52%, an improvement from the 86.65%, approaching the baseline accuracy (94.05%).
Loss Reduction: The model demonstrated steady loss reduction throughout the training process, beginning at 2.3028 during Epoch 1 in Trial 0 and decreasing to 0.5606 by Epoch 10 in Trial 1, indicating effective error minimization. This consistent decline reflects the model’s ability to optimize its parameters and improve performance across trials. One of the standout results of this experiment was the final reduction in inference time by 15.20%, dropping to 0.001043 seconds/image, marking a considerable improvement compared to the baseline DenseNet’s 0.00123 seconds. See Table 3.
A.3 ResNet-18:
@@
Loss Reduction: The loss reduction across trials also followed a clear downward trend. In Trial 3, with 0.4805 layer sparsity and a rank of 10, the validation loss dropped from 1.9847 to 0.6321 (first to tenth epoch), indicating better model convergence. Notably, inference time dropped by 36.4% to 0.00007 seconds/image, improving from the baseline of 0.00011 seconds/image See Table 4.
@@ 18. Answer: [Yes]
  19. Justification: The code and instructions on how to reproduce the experiments are included in the GitHub link. Within the paper, the model is described in detail with code snippets and is reproducible using the information provided.
@@ 23. Answer: [Yes]
  24. Justification: Code is provided with instructions through the GitHub link.